OpenStudy (anonymous):

divide 12x-12 over x-1

OpenStudy (lgbasallote):

\[\frac{12x - 12}{x-1}\] factor out the numerator \[\implies \frac{12(x-1)}{x-1}\]

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

yes thanks

OpenStudy (lgbasallote):

welcome

OpenStudy (valpey):

Make a note that when you factor something out of a fraction, you have to also exclude any cases where that thing would have been equal to zero (it is in the denominator that it is a problem because division is not defined for zero). Therefore, you can factor out x-1, but x=1 can no longer be considered as a possible value for x.